OpenStudy (dls):

Derivtative of sin^2 (x)^5? Confused..help please..

OpenStudy (abb0t):

is this correct? \[\sin^2(x^5)\]

OpenStudy (abb0t):

or \[[\sin^2(x)]^5\]

OpenStudy (dls):

first one

OpenStudy (abb0t):

You use chain rule. So derivative of the outside times derivative of the inside.

OpenStudy (dls):

\[\LARGE \cos^2(x^5) \times 2\cos(x)^5 \times 5x^4\] this is wrong? :?

OpenStudy (dls):

first sine...then sine(square) then x^5

OpenStudy (abb0t):

Your outside would be \[\sin^2(u)\] where: \[u=x^5\]

OpenStudy (abb0t):

\[2\sin(x^5) \times \cos(x^5) \times 5x^4\]

OpenStudy (abb0t):

Notice how I went from outside to inside? It's like saying \[[\sin(x)]^2\]

OpenStudy (abb0t):

my next step was to take the derivative of the inside, which is INSIDE the brackets. Which was the sine\[\sin^2([x^5])\] then I went MORE inside after i took the derivative: \[x^5\] Do you see the pattern?

OpenStudy (dls):

why did we differentiate 2 first and not sine?like i did?i tried to do the same thing.

OpenStudy (dls):

i went from outside>inside too

OpenStudy (dls):